OpenStudy (anonymous):

Helpful steps please, use the formula s=rwt to find the value of the missing variable . r= 3m s= pi/5m t=2sec.

OpenStudy (anonymous):

can I help

OpenStudy (shamim):

S=rwt Pi/5=3*w*2 w=?

OpenStudy (shamim):

U know Pi=3.1415

OpenStudy (anonymous):

I did that and got 6pi/5 and I'm not sure what I did wrong

OpenStudy (shamim):

No w=pi/(5*3*2)=?

OpenStudy (anonymous):

oh wow its just pi/30 ?

OpenStudy (shamim):

Ya. U r right

OpenStudy (anonymous):

thanks I see what I did now!

OpenStudy (shamim):

U r most welcome